OpenStudy (jack1):

Hiya, hlp plz? -12x^3 - 65x^2 - 117x - 70 = 0 whats the easiest way to solve the cubic(by hand, not calc)?

OpenStudy (shamil98):

you can factor out the negative 1 as a start

OpenStudy (jack1):

ok, so \[-1 (12x^3 + 65x^2 + 117x + 70) = 0\] ?

OpenStudy (anonymous):

Could use synthetic division to check for a factorization.

OpenStudy (jack1):

synthetic... how do i do that plz?

OpenStudy (anonymous):

After two tries, I found a factor and was able to factor the remaining trinomial by inspection. There are three rational roots.

OpenStudy (jack1):

yep, but is there an easy way to do it rather than guess an check?

OpenStudy (anonymous):

It isn't really guess and check. If there is an integer root, it has to be a factor of 70. So test to see if (x-a) divides the polynomial evenly, where a is a factor of 70. I recommend starting with the small ones first.

OpenStudy (jack1):

ok, so -2 is one of the roots so to test: (x-a) = (-12 - 2) = -14 so divide polynomial by -14 to test...?

OpenStudy (anonymous):

\[-12x^3 - 65x^2 - 117x - 70 = 0 \implies \]\[12x^3+65x^2 + 117x+ 70 = 0=(x+2)(4x+7)(3x+5) \implies\]\[x \in \left\{ -2,\frac{-7}{4},\frac{-5}{3} \right\}\]

OpenStudy (jack1):

12x3 + 65x2 + 117x + 70 = 0 =>> (x+2)(4x+7)(3x+5) its this step here that gets me, how do u do this, pen and paper, in a simple way?

OpenStudy (jack1):

@AnimalAin ...? i realise it's pretty easy to do with a calculator, but how do u solve without...?

ganeshie8 (ganeshie8):

"rational root thm/decartes/synthetic" combo gives x = -2 as one root

ganeshie8 (ganeshie8):

fine wid that @Jack1 .. or u want to solve it in general ?

OpenStudy (jack1):

just googling rational root theorum,

OpenStudy (jack1):

ok, yep i remember that

ganeshie8 (ganeshie8):

oh no, its too easy for u to suck it all in one line : rational root thm says this : ``` rational roots of a polynomial are of form : p/q p = factors of constant term q = factors of leading coefficient ```

ganeshie8 (ganeshie8):

so our guess and check limits to checking \(\pm 1, \pm 2, \pm 5, \pm 7 .... \)

OpenStudy (jack1):

descartes u lost me on and synthetic ive nerver heard of, sorry

OpenStudy (jack1):

ok, yep, i agree with 1,2,5,and 7

ganeshie8 (ganeshie8):

decartes says this : number of positive roots = number of sign changes in P(x) number of negative roots = number of sign changes in P(-x)

ganeshie8 (ganeshie8):

^^above are upper limits, complex roots can kick in too

ganeshie8 (ganeshie8):

ur polynomial : -12x^3 - 65x^2 - 117x - 70 = 0

OpenStudy (jack1):

ugh, h8 complex numbers >:(

ganeshie8 (ganeshie8):

there are no sign changes in P(x), so there will not be any positive roots. so dont bother about them.

ganeshie8 (ganeshie8):

Having the knowledge that there will not be any positive roots, out guess and check work reduces by half : \(-1, -2, -5, -7,....\)

OpenStudy (jack1):

ok, think i get descartes now too, cheers @ganeshie8 ;)

ganeshie8 (ganeshie8):

If you get lucky with any one root, then you're done. u can divide it and solve the remaining depressed quadratic using quadratic formula

OpenStudy (jack1):

so always hit the easy possible numbers first then and hope to get lucky? (1,2,3 ? )

ganeshie8 (ganeshie8):

yes, thats the easy way if it works

ganeshie8 (ganeshie8):

if u dont get lucky we can solve it analytically

OpenStudy (jack1):

sweet... cheers man while i've got u tho, could u give me a run down on synthetic division if u have a sec?

ganeshie8 (ganeshie8):

we can always solve cubic and quartic equations but we cannot solve 5th degree and above always

ganeshie8 (ganeshie8):

yeah sure :)

ganeshie8 (ganeshie8):

synthetic is just a shortcut for long division

ganeshie8 (ganeshie8):

easy example first : divide x^2 + 2x + 1 by x+1

ganeshie8 (ganeshie8):

\( \begin{array}{} -1~|&1&2&1&\\ &&&&\\ \hline &&&&\\ &1&&& \end{array} \)

ganeshie8 (ganeshie8):

1, 2, 1 are the coefficients of polynomial

ganeshie8 (ganeshie8):

since you want to divide by x+1, put x = -1 on left side

ganeshie8 (ganeshie8):

and to start with, drop that "1" vertically down and put it in the last row

ganeshie8 (ganeshie8):

thats the start...

OpenStudy (jack1):

ok, follow what you did there, will have to try a few before it becomes muscle memory tho cheers again for all this @ganeshie8 !! ur an awesome teacher / tutor hey

ganeshie8 (ganeshie8):

step 2 : 1(-1) = -1 : \( \begin{array}{} -1~|&1&~~2&1&\\ &&-1&&\\ \hline &&&&\\ &1&&& \end{array} \)

ganeshie8 (ganeshie8):

step3 : add them vertically and put the result down in the last row : \( \begin{array}{} -1~|&1&~~2&1&\\ &&-1&&\\ \hline &&&&\\ &1&~~1&& \end{array} \)

ganeshie8 (ganeshie8):

repeat step2 and step3 till u hit the end

ganeshie8 (ganeshie8):

\( \begin{array}{} -1~|&1&~~2&~~1&\\ &&-1&-1&\\ \hline &&&&\\ &1&~~1&~~|0& \end{array} \)

ganeshie8 (ganeshie8):

look at last row : 1x + 1 is ur quotient 0 is ur remainder

OpenStudy (jack1):

ahhhhh!!!! gotcha now man

ganeshie8 (ganeshie8):

so, (x^2 + 2x + 1 )/(x+1) = 1x + 1 with a remainder 0

ganeshie8 (ganeshie8):

:) you can try dividing ur polyomial by -2 now maybe

ganeshie8 (ganeshie8):

Since you know that "-2" is one root of cubic : -12x^3 - 65x^2 - 117x - 70 = 0 divide -2 and get the depressed quadratic equation

ganeshie8 (ganeshie8):

\( \begin{array}{} -2~|&-12&-65&-117&-70\\ &&&&\\ \hline &&&&\\ &-12&&& \end{array} \)

ganeshie8 (ganeshie8):

multiply (-12)(-2) and put it in second row : \( \begin{array}{} -2~|&-12&-65&-117&-70\\ &&~~~24&&\\ \hline &&&&\\ &-12&&& \end{array} \)

OpenStudy (jack1):

-12x^3 - 65x^2 - 117x - 70 = 0 so: -2 | -12 -65 -117 -70 24 ----------------------- umm..-12 -89 ???

ganeshie8 (ganeshie8):

add them vertically repeat...

ganeshie8 (ganeshie8):

u need to add, not subtract

ganeshie8 (ganeshie8):

multiply (-12)(-2) and put it in second row : \( \begin{array}{} -2~|&-12&-65&-117&-70\\ &&~~~24&&\\ \hline &&&&\\ &-12&-41&& \end{array} \)

ganeshie8 (ganeshie8):

^^

OpenStudy (jack1):

-12x^3 - 65x^2 - 117x - 70 = 0 so: -2 | -12 -65 -117 -70 24 82 -70 ----------------------- umm..-12 -41? 35 |0 is this right?

OpenStudy (jack1):

no wait

OpenStudy (jack1):

-12x^3 - 65x^2 - 117x - 70 = 0 so: -2 | -12 -65 -117 -70 24 82 70 ----------------------- umm..-12 -41? -35 |0 is this right?

ganeshie8 (ganeshie8):

yep ! so 0 is the denominator, and the depressed quadratic is : -12x^2 - 41x - 35

ganeshie8 (ganeshie8):

last row gives u `coefficients for the depressed polynomial` and the `remainder`

OpenStudy (jack1):

gotcha, so then we jam that into the quadratic equation an BAM, the other 2 roots?

ganeshie8 (ganeshie8):

that sit ! :) k gtg cya :)

OpenStudy (jack1):

cheers man, bye!